OpenStudy (anonymous):

Factor completely 3x^3 + 12x^2 + 18x. x^2 + 4x + 6 3(x^3 + 4x^2 + 6x) 3x(x^2 + 4x + 6) x(3x^2 + 12x + 18)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

GCFs... 3, and x?

OpenStudy (anonymous):

so 3x(x^2 + 9x^1 + 18)?

ganeshie8 (ganeshie8):

you're right about GCFs

ganeshie8 (ganeshie8):

check ur second step again, 9 and 18 are not correct.

OpenStudy (anonymous):

3x(x^2 + 9x^1 + 15)?

ganeshie8 (ganeshie8):

3x^3 + 12x^2 + 18x factor out GCF 3x 3x(x^3 + 4x + 6)

ganeshie8 (ganeshie8):

12/3 = 4 18/3 = 6 ok ?

OpenStudy (anonymous):

oh my bad i was subtracting lol

OpenStudy (anonymous):

3x(x^2 + 4x^1 + 6)

OpenStudy (anonymous):

now that weird formula lol

OpenStudy (anonymous):

(a+b)2=a2+2ab+b2 (a−b)2=a2−2ab+b2 a2−b2=(a+b)(a−b)

ganeshie8 (ganeshie8):

knw what, we're done ! look at ur options :)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

thanks :D

ganeshie8 (ganeshie8):

yw (: